OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below: If x = 3 and 3 = y, then x = y.

OpenStudy (anonymous):

Commutative Property of Addition Associative Property of Addition Transitive Property Symmetric Property

OpenStudy (perl):

Commutative Property of Addition example : 1 + 2 = 2 + 1 Associative Property of Addition example: 1 + ( 2 + 3) = ( 1 + 2) + 3 Transitive Property if a = b and b = c, then a = c Symmetric Property if a = b then b = a

OpenStudy (texaschic101):

wrong answer

OpenStudy (anonymous):

thank you @perl

OpenStudy (anonymous):

i go with @perl

OpenStudy (anonymous):

k im pretty sure she can figure it out on her own by what pearl said...